# A predictive modular approach to constraint satisfaction under uncertainty - with application to glycosylation in continuous monoclonal antibody biosimilar production

### v3

@@ 1 Introduction
There exist a number of methods to handle constraints in the presence of uncertainty. In standard Model Predictive Control (MPC), e.g., [10], a fixed nominal model is employed for constrained optimization and uncertainty is dealt with implicitly by adding output feedback. If some prior knowledge on the uncertainty is available, e.g., that it belongs to a given stochastic distribution or some bounded set, stochastic or robust MPC can be employed [6, 5, 42, 69, 45, 36, 7, 41]. When an a priori model of the system dynamics is lacking, model-free methods that give priority to safety and robustness can be applied, such as safe reinforcement learning [16, 22]. A typical case in process control is that a nominal dynamic model is available, but the uncertainty is not well characterized and this causes problems for methods that do not adapt to compensate for observed uncertainty during runtime, such as MPC. In such cases one can improve constraint handling by employing learning and adaptive based methods [21, 12, 61, 8, 43, 13]. However, essentially all such methods rely on particular controller structures, such as different variants of MPC, and are therefore not applicable to other control algorithms. Predictive filters [59, 63], on the other hand, are modular and can be combined with essentially any control algorithm but are focused on satisfying hard safety constraints. Predictive filters have recently been considered also for dealing with soft constraints, but then primarily to avoid infeasibilities [64, 65]. There also exist works on robust extensions of the predictive safety filter, e.g., [66], but these are all based on prior knowledge of the uncertainty. We note that constraints also can be handled with other methods, such as control barrier functions [4] and Hamilton–Jacobi reachability analysis [20]. The interested reader is referred to [62] and [11] for an overview of constraint handling methods. Finally, in the context of learning-based control, there exist works focusing on parameter adaptation with economic considerations [29, 31, 23, 30, 2, 70, 46]. However, rather than high-cost constraint satisfaction for arbitrary systems, they consider obtaining operational conditions to optimize economical performance using learning based methods. To our knowledge there exist no methods for hard and soft constraint satisfaction that can be applied to systems with arbitrary controllers while adapting optimal constraint margins for the soft constraints directly from data, requiring no prior knowledge on the uncertainty.
@@
Preliminary results have been presented in [67] using examples involving water tank temperature control and fed-batch penicillin fermentation optimization. Here we extend these results, present a realistic case study of glycosylation constraint satisfaction in continuous monoclonal antibody biosimilar production based on a realistic metabolic network and glycosylation models.
@@ 4.5 Guideline for parameter selection
The parameters of the proposed method can be selected as follows. The initial constraint margin $\theta_{0}$ is typically set to $\theta_{0}=0$, after which the constraint adapter learns the optimal margin from online data. The prediction horizon $N$ and sampling time $T_{s}$ in the predictive filter are chosen based on the available sampling rate and computational capacity, with larger $N$ improving prediction accuracy but increasing computational load. The step-size sequences $\{a_{j}\}$ and $\{c_{j}\}$ follow standard FDSA guidelines [56, 57] and may be tuned heuristically [61] if needed. The convergence threshold $\rho_{\text{learning}}$ determines when the margin is considered converged, where smaller values lead to slower but more precise convergence. The averaging window $n_{\text{monitor}}$ should be sufficiently large to reduce stochastic variation in the violation cost, while the detection threshold $\rho_{\text{monitor}}$ specifies when the margin should be re-learned. Smaller $\rho_{\text{monitor}}$ values increase sensitivity to changes in the underlying uncertainty.
@@ 5 Case Study: Glycosylation in Perfusion BioReactor
5.1 Glycosolation in Continuous mAb Production
Glycosylation, a post-translational modification of proteins produced by mammalian cells, is a critical quality attribute (CQA) of monoclonal antibodies (mAbs). It has significant impact on properties such as the pharmacokinetics and pharmacodynamics of mAbs [53, 35]. The glycosylation attributes of biosimilar products are required to be within given reference ranges - so-called specifications - to ensure product safety and efficacy [15, 33, 19]. If the reference ranges are exceeded, the biosimilar products needs to be rejected or recycled, or it must be justified that the biosimilar has a similar effect as the reference product for approval. In any case, violating the given reference range will induce significant economic losses.
@@ 5.4 Constraint handling using learning-based predictive filter
where $v_{u}\sim U(-0.05,0.05)$ and the other parameters are given in (18). We assume the $P_{G_{0}}$ can be predicted accurately by in-line measurements with sampling time $T_{s}=5\text{ min}$.
@@ 5.4.2 Results
Figure 8: Constraint adaptation and violation cost monitoring results using the learning-based predictive filter. The figure above shows the constraint margin $\theta$ value in the predictive filter, while the figure below shows the corresponding average constraint violation cost obtained by averaging (5) of 48 samples for each theta. The blue and green regions represent the $\theta$ adaptation and violation cost monitoring phases respectively.
@@
where $j$ is the number of iterations of FDSA update. The threshold for convergence is chosen to be $\rho_{learning}=1\times 10^{-3}$. For each $\theta$ value considered in the gradient approximation step, we measure the violation cost with penalty term (5) where $\lambda=10$ and $P_{k}=\max(0,P_{G_{0}}-P_{G_{0},opt})$ by averaging $48$ samples with the corresponding $\theta$ value applied in the predictive filter. As shown in Figure 8 (a), $\theta$ quickly converges to $\theta^{*}=0.026$ with a low violation cost within few iterations, starting from $\theta_{0}=0$ where the constraint violation cost is relatively high. The converged $\theta^{*}=0.026$ is in the region with low average violation cost, as illustrated in Figure 9. Note that when using the feedback controller (19) without constraint handling (blue line, Figure 8 (b)), the average constraint violation cost is $1.34$, which is more than $2.5$ times larger than the case when using the predictive filter with $\theta=\theta^{*}$ with the average constraint violation cost $0.53$. This demonstrates that constraint handling as such is important under uncertainty.
@@ 6 Conclusion and Discussion
In this paper we proposed a universal modular approach to constraint handling, based on a predictive filter combined with constraint margin adaptation and constraint violation cost monitoring to minimize the cost of soft constraint violations. The main advantage of the proposed method is that, as a universal module, it can be easily applied to a system with arbitrary controllers and can ensure constraint satisfaction with minimal modification of the desired closed-loop performance. This provides a simple approach to enable systems to complete tasks using arbitrary controllers or optimizers without significant economic cost induced by high-cost soft constraint violations. By constantly monitoring the constraint violation cost, the proposed method can adjust optimal constraint margins according to changes in the size and characteristics of uncertainties. Both the learning of the optimal constraint margin and the detection of the changes in uncertainties are directly from real-time data, and no prior knowledge of the uncertainties is required. We illustrated the potential of the proposed method using a realistic simulation case study of glycosylation constraint satisfaction in continuous mAb biosimilar production based on a realistic metabolic network and glycosylation models involving 23 extracellular metabolites and 126 reactions. The method resulted in a more than 60% reduction in the cost due to constraint violations as compared to the case without the proposed constraint handling method. The efficiency of the constraint margin adaptation and constraint satisfaction, while largely maintaining the desired closed-loop behaviour, demonstrates the potential of the proposed method for application in complex control and optimization systems that do not explicitly handle constraints.

### v4

@@ 1 Introduction
There exist a number of methods to handle constraints in the presence of uncertainty. In standard Model Predictive Control (MPC), e.g., [10], a fixed nominal model is employed for constrained optimization and uncertainty is dealt with implicitly by adding output feedback. If some prior knowledge on the uncertainty is available, e.g., that it belongs to a given stochastic distribution or some bounded set, stochastic or robust MPC can be employed [6, 5, 42, 69, 45, 36, 7, 41]. When an a priori model of the system dynamics is lacking, model-free methods that give priority to safety and robustness can be applied, such as safe reinforcement learning [16, 22]. A typical case in process control is that a nominal dynamic model is available, but the uncertainty is not well characterized and this causes problems for methods that do not adapt to compensate for observed uncertainty during runtime, such as MPC. In such cases one can improve constraint handling by employing learning and adaptive based methods [21, 12, 61, 8, 43, 13]. However, essentially all such methods rely on particular controller structures, such as different variants of MPC, and are therefore not applicable to other control algorithms. Predictive filters [59, 63], on the other hand, are modular and can be combined with essentially any control algorithm but are focused on satisfying hard safety constraints. Predictive filters have recently been considered also for dealing with soft constraints, but then primarily to avoid infeasibilities [64, 65]. There also exist works on robust extensions of the predictive safety filter, e.g., [66], but these are all based on prior knowledge of the uncertainty. We note that constraints also can be handled with other methods, such as control barrier functions [4] and Hamilton–Jacobi reachability analysis [20]. The interested reader is referred to [62] and [11] for an overview of constraint handling methods. Finally, in the context of learning-based control, there exist works focusing on parameter adaptation with economic considerations [29, 31, 23, 30, 2, 70, 46]. However, rather than high-cost constraint satisfaction for arbitrary systems, they consider obtaining operational conditions to optimize economical performance using learning based methods. To our knowledge there exist no methods for constraint satisfaction that can be applied to systems with arbitrary controllers while adapting optimal constraint margins for the soft constraints directly from data, requiring no prior knowledge on the uncertainty.
@@
Preliminary results have been presented in [67] using examples involving water tank temperature control and fed-batch penicillin fermentation optimization. Here we extend these results and present a realistic case study of glycosylation constraint satisfaction in continuous monoclonal antibody biosimilar production based on a realistic metabolic network and glycosylation models.
@@ 4.5 Guideline for parameter selection
The parameters of the proposed method can be selected as follows. The initial constraint margin $\theta_{0}$ is typically set to $\theta_{0}=0$, after which the constraint adaptor learns the optimal margin from online data. The prediction horizon $N$ and sampling time $T_{s}$ in the predictive filter are chosen based on the available sampling rate and computational capacity, with larger $N$ improving prediction accuracy but increasing computational load. The step-size sequences $\{a_{j}\}$ and $\{c_{j}\}$ follow standard FDSA guidelines [56, 57] and may be tuned heuristically [61] if needed. The convergence threshold $\rho_{\text{learning}}$ determines when the margin is considered converged, where smaller values lead to slower but more precise convergence. The averaging window $n_{\text{monitor}}$ should be sufficiently large to reduce stochastic variation in the violation cost, while the detection threshold $\rho_{\text{monitor}}$ specifies when the margin should be re-learned. Smaller $\rho_{\text{monitor}}$ values increase sensitivity to changes in the underlying uncertainty.
@@ 5 Case Study: Glycosylation in Perfusion BioReactor
5.1 Glycosylation in Continuous mAb Production
Glycosylation, a post-translational modification of proteins produced by mammalian cells, is a critical quality attribute (CQA) of monoclonal antibodies (mAbs). It has significant impact on properties such as the pharmacokinetics and pharmacodynamics of mAbs [53, 35]. The glycosylation attributes of biosimilar products are required to be within given reference ranges — so-called specifications — to ensure product safety and efficacy [15, 33, 19]. If the reference ranges are exceeded, the biosimilar products needs to be rejected or recycled, or it must be justified that the biosimilar has a similar effect as the reference product for approval. In any case, violating the given reference range will induce significant economic losses.
@@ 5.4 Constraint handling using learning-based predictive filter
where $v_{u}\sim U(-0.05,0.05)$ and the other parameters are given in (18). We assume the $P_{G_{0}}$ can be estimated accurately by in-line measurements with sampling time $T_{s}=5\text{ min}$.
@@ 5.4.2 Results
Figure 8: Constraint adaptation and violation cost monitoring results using the learning-based predictive filter. The figure above shows the constraint margin $\theta$ value in the predictive filter, while the figure below shows the corresponding average constraint violation cost obtained by averaging (5) of 48 samples for each $\theta$. The blue and green regions represent the $\theta$ adaptation and violation cost monitoring phases respectively.
@@
where $j$ is the number of iterations of FDSA update. The threshold for convergence is chosen to be $\rho_{learning}=1\times 10^{-3}$. For each $\theta$ value considered in the gradient approximation step, we measure the violation cost with penalty term (5) where $\lambda=10$ and $P_{k}=\max(0,P_{G_{0},opt}-P_{G_{0}})$ by averaging $48$ samples with the corresponding $\theta$ value applied in the predictive filter. As shown in Figure 8 (a), $\theta$ quickly converges to $\theta^{*}=0.026$ with a low violation cost within few iterations, starting from $\theta_{0}=0$ where the constraint violation cost is relatively high. The converged $\theta^{*}=0.026$ is in the region with low average violation cost, as illustrated in Figure 9. Note that when using the feedback controller (19) without constraint handling (blue line, Figure 8 (b)), the average constraint violation cost is $1.34$, which is more than $2.5$ times larger than the case when using the predictive filter with $\theta=\theta^{*}$ with the average constraint violation cost $0.53$. This demonstrates that constraint handling as such is important under uncertainty.
@@ 6 Conclusion and Discussion
In this paper we proposed a universal modular approach to constraint handling, based on a predictive filter combined with constraint margin adaptation and constraint violation cost monitoring to minimize the cost of soft constraint violations. The main advantage of the proposed method is that, as a universal module, it can be easily applied to a system with arbitrary controllers and can ensure constraint satisfaction with minimal modification of the desired closed-loop performance. This provides a simple approach to enable systems to complete tasks using arbitrary controllers or optimizers without significant economic cost induced by high-cost soft constraint violations. By constantly monitoring the constraint violation cost, the proposed method can adjust optimal constraint margins according to changes in the size and characteristics of uncertainties. Both the learning of the optimal constraint margin and the detection of the changes in uncertainties are directly from real-time data, and no prior knowledge of the uncertainties is required. We illustrated the potential of the proposed method using a realistic simulation case study of glycosylation constraint satisfaction in continuous mAb biosimilar production based on a realistic metabolic network and glycosylation models involving 23 extracellular metabolites and 126 reactions. The method resulted in a more than 60% reduction in the cost due to constraint violations as compared to the case without the proposed constraint handling method. The efficiency of the constraint margin adaptation and constraint satisfaction, while largely maintaining the desired closed-loop behavior, demonstrates the potential of the proposed method for application in complex control and optimization systems that do not explicitly handle constraints.